OpenStudy (nali):

If the internal energy of a thermodynamic system is increased by 300J while 75J of expansion work is done...? how much heat was transferred and in which direction to or from the system?

OpenStudy (kainui):

U=q+w So this says internal energy is the sum of heat (q) and work (w). Now as you do expansion work, you are losing energy by pushing against your environment so you know that the system is losing 75 joules of work. Now you can plug those in and solve for your heat. The sign on your heat represents which direction it flowed. Positive means the internal energy rose, so the heat was transferred to the system, negative means it left the system, so energy left the system by heat if that was the case.

OpenStudy (nali):

is it like the follwoing 300J + -75J = -225

OpenStudy (nali):

i put the negative because it is exothermic

OpenStudy (kainui):

It looks more like U=q+w 300=q-75 You've set up the equation incorrectly. It's not exothermic or endothermic because there's no reaction going on here.

OpenStudy (kainui):

Actually I take that last part back about the reaction taking place.

OpenStudy (nali):

oh ok

OpenStudy (kainui):

Exothermic and endothermic just indicate whether or not heat is leaving or entering the system. So in this case, since q ends up being 375, which is positive, we know that it's endothermic since it absorbed heat.

OpenStudy (nali):

ok then thanku :D